OpenStudy (anonymous):

Find the value of x and y in this system of equations. x^y=1/y^2 and y^x=1/sqrt(x).

OpenStudy (anonymous):

\[x ^{y}=1\div y ^{2} and y ^{x}=1/\sqrt{x} \] u mean??

OpenStudy (anonymous):

From the second one, can I say that y^2 = 1/sqrt2...

OpenStudy (anonymous):

ok. Keep going.

OpenStudy (anonymous):

It´s right.

OpenStudy (anonymous):

So that gives you y and you can sub in the first to get x...

OpenStudy (anonymous):

I solved this question before, but i can´t remember all the steps.

OpenStudy (anonymous):

May you show me, how?

OpenStudy (ash2326):

@viniterranova Let's continue, would you mind showing me your work?

OpenStudy (ujjwal):

From the second i can say that y=1

OpenStudy (anonymous):

I mad x^y = y.

OpenStudy (anonymous):

So, if x^y=y I got the following expression. y= 1/y^2.

OpenStudy (ujjwal):

I don't know how you got these.. But i am sure they (your above expressions) will still give correct answer since i get y=1 from 2nd expression!!

OpenStudy (ash2326):

I know you are substituting but don't use the same variable \[x^y=y\] This will be troublesome, use another variable, if you want to but I don't see how it'd help @viniterranova

OpenStudy (anonymous):

Thanks ash2326.

OpenStudy (anonymous):

If I mad x^y=y, and rewriting the expressoin, i can got y-1/y^2=0. What do they think?

OpenStudy (ash2326):

How could you use the same variable "y"? it's defined differently. I have something. Let me show you

OpenStudy (anonymous):

ok.

OpenStudy (ash2326):

\[x^y=\frac{1}{y^2}\] \[y^x=\frac{1}{\sqrt x}\] Take log both sides, for both the equations \[y\log x=-2\log y\] \[x\log y=-\frac 1 2\log x\] multiply both the equations, we would get \[xy\log x \times \log y=\log x \times\log y\] or \[xy=1\] We have the equation \[y\log x=-2\log y\] put x=1/y \[y\log{\frac 1y}=-2\log y\] we get \[y\times ( -\log y)=-2\log y\] \[y=2\] so \[x=\frac 12\]

OpenStudy (ash2326):

@viniterranova do you get this?

OpenStudy (anonymous):

\[y\times ( -\log y)=-2\log y\]gives y=1 and y=2 ------

OpenStudy (ash2326):

Yes @mukushla I missed that then y=1 and x=1

OpenStudy (anonymous):

neat sol

OpenStudy (anonymous):

Great.

OpenStudy (anonymous):

It´s rigth.

OpenStudy (anonymous):

I got y=1 and x= 1 in another way.

OpenStudy (anonymous):

i think x=1/2 and y=2 from xy=1 if x=1/2 then y/2=1 y=2 ans

OpenStudy (anonymous):

from y log x =-2log y sub x=1/y y log(1/y)=-2log y y(-log y)=-2log y dividing both sides by -log y gives y=-2log y/(-log y) y=2 ans..... and from xy=1 2x=1 then x=1/2 ans ....

OpenStudy (anonymous):

Thanks everyone for the help.

OpenStudy (anonymous):

ok yw, have fun....